fireworks are shot upward with an initial velocity of 128 feet per second from a platform that is 3 feet above the ground. What is the peak of the fireworks? When does it get to the peak.
Answer by Alan3354(69443)

fireworks are shot upward with an initial velocity of 128 feet per second from a platform that is 3 feet above the ground. What is the peak of the fireworks? When does it get to the peak.
=================
Using 32 ft/sec/sec for gravity:
h(t) = -16t^2 + 128t + 3
---
The peak is the vertex of the parabola at t = -b/2a
t = -128/(-32) = 4 seconds
======================
Peak = h(4) = -16*4^2 + 128*4 + 3
Peak = 259 feet